Company ABC has the following capital structure:
Sh
Ordinary share capital 300000
Retained earnings 150000
8% preference share capital 100000
10% debt finance 50000
Total Capital employed 600000
This company has been paying ordinary dividend of 10% per annum and expects equity to grow
at 6% per annum. Tax = 50%
Calculate the weighted Average Cost of Capital for this com�pany
1
Expert's answer
2020-08-23T17:06:56-0400

we know the formula of weighted average cost of capital ( WACC ):

WACC = ("\frac{E}{V}" x Re ) + [ ("\frac{D}{V}" X Rd ) (1-Tc) ]

Where :

E=Market value of the firm’s equity=300000+150000=450000

D=Market value of the firm’s debt=100000+50000=150000

V=E+D=-450000+150000=600000 = total capital employed

Re=Cost of equity = 6%

Rd=Cost of debt = 10%

Tc=Corporate tax rate = 50%

WACC = ("\frac{450000}{600000}" x "\frac{6}{100}" ) + [ ("\frac{150000}{600000}" x "\frac{10}{100}" ) (1 - "\frac{50}{100}" ) ]


= ("\frac{45}{60}" x "\frac{6}{100}" ) + [ ("\frac{15}{60}" x "\frac{1}{10}" )( 1 - "\frac{1}{2}" )

= ( 0.045) + [ (0.025)(0.5)]

= (0.045) + (0.0125)

= 0.0575

WACC = 5.75%

5.75% is  the weighted Average Cost of Capital for this company.